OpenStudy (anonymous):

EASY MEDAL AND FAN What is the frequency of a microwave that has a wavelength of 2.5 cm

OpenStudy (anonymous):

What have you done so far with the problem?

OpenStudy (anonymous):

3.00x 10^8= f(25)

OpenStudy (anonymous):

answer is 1.20 x 10^10 hz

OpenStudy (anonymous):

Okay, So do you think it's correct? What method did you use?

OpenStudy (anonymous):

i think its wrong

OpenStudy (anonymous):

convert wavelength to m first, then apple the formula: \(\sf \Large f= \frac{c}{\lambda}=\frac{3 \times 10^8m/s}{0.025m}\) what will you get?